Kathy bicycles 4/km/h faster than Robert. In the same time it takes Robert to bicycle 57 km, Kathy can bicycle 69km. How fast does each bicyclist travel?
Robert -
Kathy -
Thank you,
Diane

Answer by scott8148(6628)

"Kathy bicycles 4/km/h faster than Robert" ___ K = R + 4

"In the same time it takes Robert to bicycle 57 km, Kathy can bicycle 69km" ___ t = d / r
___ the times are the same (equal) ___ (57 / R) = (69 / K)

"cross" multiplying ___ 57K = 69R

substituting ___ 57R + 228 = 69R

solve for R and substitute back to find K